Please help me with another problem:
according to the local zoo keeper,an alligator's head is about 1/3 the lenght of i'ts tail.Its torso is about 5/6 the lenght of its tail.
*gator's tail is about four feet long. how long is gator

*if the head lenght of gator's friend is about one foot long,howlong is her tail? what is her entire lenght?
*another friend has a head that is about 1 1/2 feet long,how long is her tail? what is her entire lenght?

Answer by stanbon(75887)

according to the local zoo keeper,an alligator's head is about 1/3 the length of its tail.Its torso is about 5/6 the length of its tail.
Let the length of the tail be "x"
Length of the head = 1/3 x
Length of the torso is 5/6 x
------------------

*gator's tail is about four feet long. how long is gator
EQUATION' x= 4 ft
gator = x + (1/3)x + (5/6)x
gator = 4 + (4/3) + 20/6
gator = 24/6 + 8/6 + 20/6 = 52/6 = 8 2/3 ft
=================
*if the head lenght of gator's friend is about one foot long,how
long is her tail?
EQUATION: (1/3)x = 1 ft.
x = 3 ft (length of tail)
------------
what is her entire lenght?
gator = x+ 1/3x + 5/6 x
gator = 3 + (1/3)3 + (5/6)3
gator = 3 + 1 + 2.5
gator = 6.5 ft
===========================

*another friend has a head that is about 1 1/2 feet long,how long is her tail? what is her entire length?
Comment: I'll leave this one to you.
Cheers,
Stan H.